OpenStudy (anonymous):

What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

OpenStudy (anonymous):

please explain how you got it or show work

OpenStudy (zpupster):

point-slope form y-y1= m(x-x1) so just plug in the numbers x1 = 4 and y1 =7 from (4,7) y-7 = 3(x-4) but you still should put it in this format y=mx+b

OpenStudy (anonymous):

can you show me the final answer in that format please

OpenStudy (anonymous):

is it y=3x-5

OpenStudy (zpupster):

y-7 = 3(x-4) first mult both terms in parentheses y-7 = 3x-12 then add 7 to both sides y= 3x-5

OpenStudy (zpupster):

see the point on the line